OpenStudy (nuccioreggie):

The table below shows the number of cookies in different numbers of packs: Number of Packs Number of Cookies 2 6 3 9 4 12 5 25 Is this true or false? The numbers in the table represent a proportional relationship. True False

OpenStudy (tkhunny):

Divide Cookies by Packs. See if you get the same result each time.

OpenStudy (nuccioreggie):

I did so that mean its true?? @tkhunny

OpenStudy (tkhunny):

6/2 = 3 9/3 = 3 12/4 = 3 25/5 = 5 -- Whoops! Did you mean 15?

OpenStudy (nuccioreggie):

False